Suppose that Y increases $1,000 after G increases $500. What is MPC?
Answer #1

Change in Y=K*Change in G

Change in Y=1000

Change in G=500

K=1000/500

K=2

K=1/1-MPC

2=1/1-MPC

MPC=.5